OpenStudy (anonymous):

what is the solution to the (3*10^2)*(5*10^5) written in scietific notation 1.5*10^8 15*10^8 1.5*10^7 15*10^7

OpenStudy (anonymous):

You should add your exponents first and then multiply the leading integers, and then put the numbers into scientific notation (x*10^y) where x is less than 10.

OpenStudy (anonymous):

I'll help you \[(3\times10^{2}) \times (5\times10^{5}) = (3\times5)\times(10^{2+5}) = 15\times10^{7} = 1.5 \times10^{8}\]

OpenStudy (anonymous):

thank you sooooooooooooo much